Morgan takes 3 hours to paint a room and Jenae take 5 hours. How long will it take them working together?
Answer by unlockmath(1688)

Hello,
We can set this up as:
x/3+x/5=1
Multiply by 15:
5x+3x=15
Combine and divide by 5 to get:
x=15/8 or 1 7/8 hours
Make sense?
RJ
www.math-unlock.com